# TrustChain: A Blockchain Framework for Auditing and Verifying Aggregators in Decentralized Federated Learning

### v1

@@ VI-B Results Analysis
From Fig. 2(g–i), it can be inferred that the portion of Byzantine nodes has a greater impact on their potency compared to the case when TrustChain is used. However, this observation is less noticeable for label flipping, as it varies within a smaller range. Moreover, TOM appears to be more difficult to mitigate, as it results in the lowest classification accuracy in most experiments shown in Fig. 2. For large numbers of Byzantine nodes, sign flipping can also lead to catastrophic errors, when TrustChain is not employed. Nevertheless, the results indicate that TrustChain can mitigate the effect of the Byzantine population to a large extent. This can be seen by comparing panels (d–f) and (g–i) in Fig. 2.
Abblation study
@@
Significance test

### v2

@@ VI-B2 Sensitivity to attacks
From Fig. 2, it can be inferred that the portion of Byzantine nodes has a greater impact on their potency compared to the case when TrustChain is used. However, this observation is less noticeable for label flipping, as it varies within a smaller range. Moreover, TOM appears to be more difficult to mitigate, as it results in the lowest classification accuracy in most experiments shown in Fig. 2. For large numbers of Byzantine nodes, sign flipping can also lead to catastrophic errors, when TrustChain is not employed. Nevertheless, the results indicate that TrustChain can mitigate the effect of the Byzantine population to a large extent. This can be seen by comparing plots in the middle and right side of Fig. 2.
VI-B3 Ablation study
@@
The overall success rates of PSE and PAA for each attack type are presented in Fig. 4. The consistently high detection rates of these two components demonstrate the effectiveness of SC1 and SC2 in mitigating malicious actors both before and after aggregator selection in each round. While the success rate remains high across all attacks, TOM proves more challenging to detect, resulting in a slightly lower success rate compared to the other attack types.
Figure 4: Overall detection performance of PSE (used in SC1) and PAA (used SC2) for different attacks.
VI-B4 Significance test